(1 point) The probability of the closing of the ith relay in the circuits shown is given by Pi. Let P0.4, P2 that a current flows between A and B for the respective circuits? 0.7, Ps0.5, p-0.3,p0.2.f all relays function independently. what is the probability (a) P (b) P-
Answer #1

a)for above to be worked: P= (1-12 not work*34 not work)*5 works =(1-(1-0.4*0.7)*(1-0.5*0.3))*0.2=0.0776

b)in same manner probability that current flows =p1p4+p1p3p5+p2p5+p2p3p4-p1p3p4p5-p1p2p4p5-p1p2p3p4-p1p2p3p5-p2p3p4p5+p1p2p3p4p5 =0.2936